# Categorical Reparameterization with Denoising Diffusion models

### v1

@@ B.1 An alternative view of ReinMax
The first term equals $\tfrac{1}{2}(\Sigma_{1}+\Sigma_{2})$. For the second term,
$\mathbb{C}\mathrm{ov}\big(\mathbb{E}[X\mid Z]\big)=\mathbb{C}\mathrm{ov}(\mu_{Z})=\mathbb{E}[\mu_{Z}\mu_{Z}^{\top}]-\mu\mu^{\top}=\tfrac{1}{2}(\mu_{1}\mu_{1}^{\top}+\mu_{2}\mu_{2}^{\top})-\mu\mu^{\top}.$
With $\mu=\frac{\mu_{1}+\mu_{2}}{2}$ a short algebraic simplification gives
$\frac{1}{2}(\mu_{1}\mu_{1}^{\top}+\mu_{2}\mu_{2}^{\top})-\mu\mu^{\top}=\frac{1}{4}(\mu_{1}-\mu_{2})(\mu_{1}-\mu_{2})^{\top},$
and combining the two terms yields the stated formula. ∎
B.2 Exactness of ReinMax for quadratic functions
In this section we show that (37) is unbiased for quadratic functions. Let $f(x)=x^{\top}Ax+b^{\top}x+d,$ with $A$ a symmetric matrix. Then $\nabla_{\!{x}}f(x)=2Ax+b$ is affine. Moreover,
$\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\mathbb{C}\mathrm{ov}(\pi_{\!\theta})\,\overline{f},$
where $\overline{f}=(f(e_{i}),\dots,f(e_{K}))^{\top}$ and we used the identity $\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}=\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)$. Since $f(e_{i})=A_{ii}+b_{i}+d$, denoting $\overline{f}=\mathrm{diag}(A)+b+d\mathbf{1}$ where $\mathrm{diag}(A)\in{\mathbb{R}}^{K}$ is the vector of diagonal entries of $A$,
$\nabla_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\big(\mathrm{diag}(A)+b\big),$ (38)
where we have used that $\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\mathbf{1}=0$. We now compute exactly $\mathbb{E}_{\pi_{\!\theta}}[\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)]$. Write $C_{\theta}:=\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)$ and $\mu_{\theta}\coloneqq\mathbb{E}_{\pi_{\!\theta}}[X]$. Using the definition (37),
@@ B.2 Exactness of ReinMax for quadratic functions
$\mathbb{E}_{\pi_{\!\theta}}[\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)]$ $=\frac{1}{2}\Big(C_{\theta}\,\mathbb{E}_{\pi_{\!\theta}}[2AX+b]+\mathbb{E}_{\pi_{\!\theta}}\big[(X-\mu_{\theta})(X-\mu_{\theta})^{\top}(2AX+b)\big]\Big)$
$=\frac{1}{2}\Big(C_{\theta}(2A\mu_{\theta}+b)+2\,\mathbb{E}_{\pi_{\!\theta}}\big[(X-\mu_{\theta})(X-\mu_{\theta})^{\top}AX\big]+C_{\theta}b\Big).$ (39)
It remains to evaluate $T:=\mathbb{E}_{\pi_{\!\theta}}[(X-\mu_{\theta})(X-\mu_{\theta})^{\top}AX]$. We have that
$T=\sum_{k=1}^{K}\pi_{\!\theta}(e_{k})\,(e_{k}-\mu_{\theta})(e_{k}-\mu_{\theta})^{\top}Ae_{k}.$
Using $e_{k}^{\top}Ae_{k}=A_{kk}$ and $\mu^{\top}Ae_{k}=(A\mu)_{k}$, expand
$(e_{k}-\mu)(e_{k}-\mu)^{\top}Ae_{k}=A_{kk}e_{k}-(A\mu)_{k}e_{k}-A_{kk}\mu+(A\mu)_{k}\mu.$
Summing over $k$ yields
$T=\big(\mathrm{diag}(\mu)-\mu\mu^{\top}\big)\mathrm{diag}(A)-\big(\mathrm{diag}(\mu)-\mu\mu^{\top}\big)(A\mu)=C\,\mathrm{diag}(A)-C(A\mu).$ (40)
Plugging (40) into (39) gives
@@
$\mathbb{E}_{\pi_{\!\theta}}[\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)]=\frac{1}{2}\Big(C(2A\mu+b)+2(C\mathrm{diag}(A)-C(A\mu))+Cb\Big)=C\big(\mathrm{diag}(A)+b\big)\;,$
which proves the result
@@ C.1 Reverse transitions
where we assume the more general reference $\pi_{1}=\mathcal{N}(\mu,\Sigma)$ with $\Sigma\in\mathcal{S}_{++}({\mathbb{R}}^{d})$. Let $(\eta_{t})_{t\in[0,1]}$ be a schedule such that $0\leq\eta_{t}\leq\sigma_{t}$.
From the following equality in law which follows from the fact that $\pi_{1}$ is a Gaussian distribution,
$\sigma_{t}X_{1}\overset{\mathcal{L}}{=}(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}\tilde{X}_{1}+\big(\sigma_{t}-(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}\big)\mu+\eta_{t}\Sigma^{1/2}Z\;,\quad(\tilde{X}_{1},Z)\sim\pi_{1}\otimes\mathcal{N}(0,\mathbf{I}_{d})$
we find that
$X_{t}\overset{\mathcal{L}}{=}\alpha_{t}X_{0}+(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}X_{1}+\big(\sigma_{t}-(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}\big)\mu+\eta_{t}\Sigma^{1/2}Z_{t},$ (42)
where $(X_{0},X_{1},Z_{t})\sim\pi_{0}\otimes\pi_{1}\otimes\mathcal{N}(0,\mathbf{I}_{d})$. We thus define for any $t\in[0,1]$ the following bridge transitions,
$q^{\eta}_{\smash{t\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{t}|x_{0},x_{1})\coloneqq\begin{cases}\updelta_{\alpha_{t}x_{0}+\sigma_{t}x_{1}}(\mathrm{d}x_{t}),&\eta_{t}^{2}=0\;,\\ \mathrm{N}(x_{t};\alpha_{t}x_{0}+(\sigma^{2}_{t}-\eta_{t}^{2})^{1/2}x_{1}+\big(\sigma_{t}-(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}\big)\mu,\,\eta^{2}_{t}\Sigma)\mathrm{d}x_{t},&\eta_{t}^{2}>0\;.\end{cases}$ (43)
Then clearly from (41) and (42) we have for all $\eta_{t}\in[0,\sigma_{t}]$,
$\pi_{t}(\mathrm{d}x_{t})=\int q^{\eta}_{\smash{t\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{t}|x_{0},x_{1})\,\pi_{0}(x_{0})\pi_{1}(x_{1})\mathrm{d}x_{0}\mathrm{d}x_{1}\;.$ (44)
Now define the reverse transition
@@
where $\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x_{t})$ denotes the conditional law of $(X_{0},X_{1})$ given $X_{t}=x_{t}$ under the joint distribution induced by (41). This conditional can be written as $\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})=\updelta_{(x_{t}-\alpha_{t}x_{0})/\sigma_{t}}(\mathrm{d}x_{1})\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}x_{0}|x_{t}),$
$\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})=\frac{\pi_{0}(x_{0})\,\mathrm{N}(x_{t};\alpha_{t}x_{0}+\sigma_{t}\mu,\sigma_{t}^{2}\Sigma)}{\pi_{t}(x_{t})}.$ (46)
Indeed, for any bounded measurable function $f$,
$\int f(x_{0},x_{t},x_{1})\,\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})\pi_{t}(\mathrm{d}x_{t})$ $=\int f(x_{0},x_{t},x_{1})\,\updelta_{\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}}}(\mathrm{d}x_{1})\pi_{0}(x_{0})\mathrm{N}(x_{t};\alpha_{t}x_{0}+\sigma_{t}\mu,\sigma_{t}^{2}\Sigma)\mathrm{d}x_{0}\mathrm{d}x_{t}$
$=\int f(x_{0},x_{t},\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}})\,\pi_{0}(x_{0})\mathrm{N}(x_{t};\alpha_{t}x_{0}+\sigma_{t}\mu,\sigma_{t}^{2}\Sigma)\mathrm{d}x_{0}\mathrm{d}x_{t}$
$=\int f(x_{0},\alpha_{t}x_{0}+\sigma_{t}\mu+\sigma_{t}\Sigma^{1/2}z,\mu+\Sigma^{1/2}z)\,\pi_{0}(x_{0})\mathrm{N}(z;0,\mathbf{I}_{d})\,\mathrm{d}x_{0}\mathrm{d}z$
$=\int f(x_{0},\alpha_{t}x_{0}+\sigma_{t}x_{1},x_{1})\,\pi_{0}(x_{0})\mathrm{N}(x_{1};\mu,\Sigma)\mathrm{d}x_{0}\mathrm{d}x_{1}$
$=\int f(x_{0},x_{t},x_{1})\,\updelta_{\alpha_{t}x_{0}+\sigma_{t}x_{1}}(\mathrm{d}x_{t})\pi_{0}(x_{0})\mathrm{N}(x_{1};\mu,\Sigma)\mathrm{d}x_{0}\mathrm{d}x_{1}$
@@
$\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})\pi_{t}(\mathrm{d}x_{t})=\updelta_{\alpha_{t}x_{0}+\sigma_{t}x_{1}}(\mathrm{d}x_{t})\pi_{0}(x_{0})\pi_{1}(x_{1})\mathrm{d}x_{0}\mathrm{d}x_{1}\;,$ (47)
where the r.h.s. is the joint distribution defined by (41). It then follows that
@@
$=\int q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{s}|x_{0},x_{1})\,\pi_{0}(x_{0})\pi_{1}(x_{1})\mathrm{d}x_{0}\mathrm{d}x_{1}$
$=\pi_{s}(x_{s})$
@@
$=\int\underbrace{q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{s}|x_{0},\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}})}_{q^{\eta}_{\smash{s|0,t}}(\cdot|x_{0},x_{t})}\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\mathrm{d}x_{0}$
where the defined $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(\cdot|x_{0},x_{t})$, up to the notation, is exactly the DDIM bridge transition song2021ddim when $\mu=0_{d}$ and $\Sigma=\mathbf{I}_{d}$. Finally, the Gaussian approximation $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(\cdot|\hat{x}_{0}(x_{t},t),x_{t})$ used at inference, with $\hat{x}_{0}(x_{t},t)\coloneqq\int x_{0}\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\mathrm{d}x_{0}$, is the one solving
$\operatorname*{argmin}_{r_{s|t}(\cdot|x_{t})\in\mathcal{G}_{\eta^{2}_{s}\Sigma}}\mathsf{KL}(\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x_{t})\|r_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x_{t}))\;,$

### v2

@@ B.2 ReinMax estimator
The first term is
$\mathbb{E}\big[\mathbb{C}\mathrm{ov}(X\mid Z)\big]=\tfrac{1}{2}\Sigma_{1}+\tfrac{1}{2}\Sigma_{2}.$
For the second term, since $\mathbb{E}[X\mid Z]=\mu_{Z}$ takes values $\mu_{1}$ and $\mu_{2}$,
$\mathbb{C}\mathrm{ov}(\mu_{Z})=\mathbb{E}[\mu_{Z}\mu_{Z}^{\top}]-\mu\mu^{\top}=\tfrac{1}{2}(\mu_{1}\mu_{1}^{\top}+\mu_{2}\mu_{2}^{\top})-\mu\mu^{\top}.$
Using $\mu=\frac{\mu_{1}+\mu_{2}}{2}$, we expand
$\mu\mu^{\top}=\frac{1}{4}\big(\mu_{1}\mu_{1}^{\top}+\mu_{1}\mu_{2}^{\top}+\mu_{2}\mu_{1}^{\top}+\mu_{2}\mu_{2}^{\top}\big),$
hence
$\tfrac{1}{2}(\mu_{1}\mu_{1}^{\top}+\mu_{2}\mu_{2}^{\top})-\mu\mu^{\top}$ $=\frac{1}{4}\big(\mu_{1}\mu_{1}^{\top}-\mu_{1}\mu_{2}^{\top}-\mu_{2}\mu_{1}^{\top}+\mu_{2}\mu_{2}^{\top}\big)$
$=\frac{1}{4}(\mu_{1}-\mu_{2})(\mu_{1}-\mu_{2})^{\top}.$
Combining the two terms gives the stated expression for $\Sigma$. ∎
We now consider the extension to $L>1$. The ReinMax estimator is given by
$\widehat{\nabla}^{\text{RM}}_{\!\varphi}G(X;\varphi)\coloneqq\frac{1}{2}\,B_{\varpi_{\!\varphi}}(X)\nabla_{\!{x}}f(X)\;,$ (46)
where $X\sim\varpi_{\!\varphi}$ and
$B_{\varpi_{\!\varphi}}(X)=\mathbb{C}\mathrm{ov}_{\varpi_{\!\varphi}}(X)+\widehat{C}_{\varphi}(X)$
where $\widehat{C}_{\varphi}(X)$ is block-diagonal with $L$ blocks of size $K\times K$; its $\ell$-th block is $\widehat{C}^{(\ell)}_{\varphi}(X)\coloneqq\bigl(X^{\ell}-\mathbb{E}_{\varpi_{\!\varphi^{\ell}}}[X^{\ell}]\bigr)\bigl(X^{\ell}-\mathbb{E}_{\varpi_{\!\varphi^{\ell}}}[X^{\ell}]\bigr)^{\top}$, $\ell\in[L]$, so that $\mathbb{E}_{\varpi_{\!\varphi}}[\widehat{C}^{(\ell)}_{\varphi}(X)]=\mathbb{C}\mathrm{ov}_{\varpi_{\!\varphi}}(X^{\ell})$. The matrix $B_{\varpi_{\!\varphi}}(X)$ is also block-diagonal with $L$ blocks of size $K\times K$; its $\ell$-th block is
$B^{\ell}_{\varpi_{\!\varphi^{\ell}}}(X^{\ell}):=[B_{\varpi_{\!\varphi}}(X)]^{\ell,\ell}=\mathbb{C}\mathrm{ov}_{\varpi_{\!\varphi^{\ell}}}(X^{\ell})+\bigl(X^{\ell}-\mathbb{E}_{\varpi_{\!\varphi^{\ell}}}[X^{\ell}]\bigr)\bigl(X^{\ell}-\mathbb{E}_{\varpi_{\!\varphi^{\ell}}}[X^{\ell}]\bigr)^{\top})\,,\quad\ell\in[L].$
The vector $\widehat{\nabla}^{\text{RM}}_{\!\varphi}G(X;\varphi)$ is a block vector with $L$ blocks of size $K\times 1$ blocks; the $\ell$-th block is given by
$\widehat{\nabla}^{\text{RM}}_{\!\varphi}G(X;\varphi)]^{\ell}=\frac{1}{2}B^{\ell}_{\varpi_{\!\varphi^{\ell}}}(X^{\ell})\nabla_{\!{x^{\ell}}}f(X)\;.$
Taking the expectation yields to
@@
$\mathbb{E}_{\varpi_{\!\varphi}}\left[[\widehat{\nabla}^{\text{RM}}_{\!\varphi}G(X;\varphi)]^{\ell}\right]=\frac{1}{2}\mathbb{E}_{\varpi_{\!\varphi}}[B^{\ell}_{\varpi_{\!\varphi^{\ell}}}(X^{\ell})\nabla_{\!{x^{\ell}}}f^{\ell}(X^{\ell})]$
where $f^{\ell}$ is defined in (32). (44) shows that, for $\ell\in[L]$,
$\hat{\nabla}^{\text{2nd}}_{\varphi}\left[\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\right]^{\ell}=\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi^{\ell}}}[f^{\ell}(X^{\ell})]=\mathbb{E}_{\varpi_{\!\varphi}}\left[[\widehat{\nabla}^{\text{RM}}_{\!\varphi}G(X;\varphi)]^{\ell}\right]$
B.3 ReinDGE, a ReinMax based gradient estimator
Recall, following the derivations in Section 3.3, that $h_{\theta}(x_{t_{1}})\coloneqq\sum_{x_{0}}f(x_{0})\,\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x_{0}|x_{t_{1}})$. The ReinDGE gradient estimator at $\theta=\theta^{\prime}$ is given by
$\widehat{\nabla}^{\text{RM}}_{\!\theta}h_{\theta}(x_{0};T^{\theta^{\prime}}_{{t_{1}}}(X_{t_{1}})){\big|_{\theta=\theta^{\prime}}}+\operatorname{J}_{\theta}\!T^{\theta}_{t_{1}}(X_{1})^{\top}{\big|_{\theta=\theta^{\prime}}}\!\widehat{\nabla}^{\text{ST}}_{\!x_{t_{1}}}h_{\theta}(x_{0};T^{\theta^{\prime}}_{t_{1}}(X_{1}))$ (47)
where $X_{0}\sim\pi^{\theta^{\prime}}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|T^{\theta^{\prime}}_{{t_{1}}}(X_{t_{1}}))$ and following the previous section and (43), we define for any $(x_{0},x_{t_{1}})$,
@@ B.3 ReinDGE, a ReinMax based gradient estimator
$\widehat{\nabla}^{\text{ST}}_{\!x_{t_{1}}}h_{\theta}(x_{0};x_{t_{1}})$ $\coloneqq\frac{\alpha_{t_{1}}}{2\sigma^{2}_{t_{1}}}\,\mathbb{C}\mathrm{ov}_{\pi^{\theta}_{0\mathrel{\raisebox{0.32289pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|x_{t_{1}})}(X)\nabla_{\!{x}}f(x_{0})$
$\widehat{\nabla}^{\text{RM}}_{\!\theta}h_{\theta}(x_{0};x_{t_{1}})$ $\coloneqq\frac{1}{2}\,\operatorname{J}_{\theta}\!\varphi_{\theta}^{\top}B_{\theta}(x_{0};x_{t_{1}})\nabla_{\!{x}}f(x_{0})\;,$
and $B_{\theta}(x_{0};x_{t_{1}})=\mathbb{C}\mathrm{ov}_{\pi^{\theta}_{0\mathrel{\raisebox{0.32289pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|x_{t_{1}})}(X_{0})+\widehat{C}_{\theta}(x_{0};x_{t_{1}})$, where $\widehat{C}_{\theta}(x_{0};x_{t_{1}})$ is block-diagonal with $L$ blocks of size $K\times K$; its $\ell$-th block is
$\widehat{C}^{(\ell)}_{\theta}(x_{0};x_{t_{1}})\coloneqq(x^{\ell}_{0}-\mathbb{E}_{\pi^{\theta}_{0\mathrel{\raisebox{0.32289pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|x_{t_{1}})}[X^{\ell}])(x^{\ell}_{0}-\mathbb{E}_{\pi^{\theta}_{0\mathrel{\raisebox{0.32289pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|x_{t_{1}})}[X^{\ell}])^{\top}\;.$
When using a single diffusion step (${t_{1}}=1$) and under the boundary condition $\alpha_{1}=0$, the reverse transition no longer depends on the conditioning state (equivalently $X_{1}$ carries no information about $X_{0}$), so $h_{\theta}(x_{1})$ is constant in $x_{1}$ and $h_{\theta}(x_{1})=\mathbb{E}_{\pi_{0}^{\theta}}\big[f(X_{0})\big].$ Hence $\nabla_{\!{x_{1}}}h_{\theta}(x_{1})=0$ and the second term in (47) vanishes. In that case, ReinDGE reduces to precisely the ReinMax gradient estimator for the categorical law $\pi_{\!\theta}$. Therefore, in the same way that ReDGE recovers Straight-Through as a special case, ReinDGE recovers ReinMax as ${t_{1}}\to 1$.
Since the derivation of this estimator may seem abstract, we provide a code snippet of its implementation in Fig. 11.
@@ C.1 Reverse transitions
and we take the (more general) reference distribution $\pi_{1}=\mathcal{N}(\mu,\Sigma)$ with $\Sigma\in\mathcal{S}_{++}({\mathbb{R}}^{d})$. Let $(\eta_{t})_{t\in[0,1]}$ be a schedule such that $0\leq\eta_{t}\leq\sigma_{t}$. In the sequel, whenever a distribution is absolutely continuous w.r.t. the Lebesgue measure, we use the same notation for the distribution and its p.d.f.
If $U$ and $V$ are random variables, $U\overset{\mathcal{L}}{=}V$ denotes equality in distribution, i.e. $\text{Law}(U)=\text{Law}(V)$. Since $\pi_{1}$ is Gaussian, we may decompose the Gaussian variable $\sigma_{t}X_{1}$ as
$\sigma_{t}X_{1}\overset{\mathcal{L}}{=}(\sigma_{t}^{2}-\eta_{t}^{2})^{1/2}X_{1}+\Bigl(\sigma_{t}-(\sigma_{t}^{2}-\eta_{t}^{2})^{1/2}\Bigr)\mu+\eta_{t}\Sigma^{1/2}Z,\qquad(X_{1},Z)\sim\pi_{1}\otimes\mathcal{N}(0,\mathbf{I}_{{d}})\;.$ (49)
Indeed, both sides have mean $\sigma_{t}\mu$ and covariance $\sigma_{t}^{2}\Sigma$.
Combining (48) and (49), we obtain $X_{t}\overset{\mathcal{L}}{=}X_{t}^{\eta}$, where we define
$X_{t}^{\eta}$ $=\alpha_{t}X_{0}+(\sigma_{t}^{2}-\eta_{t}^{2})^{1/2}X_{1}+\Bigl(\sigma_{t}-(\sigma_{t}^{2}-\eta_{t}^{2})^{1/2}\Bigr)\mu+\eta_{t}\Sigma^{1/2}Z_{t},$ (50)
$\hskip 56.9055pt(X_{0},X_{1},Z_{t})\sim\pi_{0}\otimes\pi_{1}\otimes\mathcal{N}(0,\mathbf{I}_{{d}})\;.$
Denote by $q^{\eta}_{\smash{t\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\cdot|x_{0},x_{1})$ the conditional distribution of $X_{t}^{\eta}$ given $(X_{0},X_{1})=(x_{0},x_{1})$. Then clearly from (48) and (50) we have for all $\eta_{t}\in[0,\sigma_{t}]$,
$\pi_{t}(\mathrm{d}x_{t})=\int q^{\eta}_{\smash{t\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{t}|x_{0},x_{1})\,\pi_{0}(\mathrm{d}x_{0})\pi_{1}(\mathrm{d}x_{1})\;.$ (51)
Now, for $0\leq s<t\leq 1$, define the reverse transition
@@
where $\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x_{t})$ denotes the conditional distribution of $(X_{0},X_{1})$ given $X_{t}=x_{t}$ under the joint distribution induced by (48). This conditional can be written as $\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})=\updelta_{(x_{t}-\alpha_{t}x_{0})/\sigma_{t}}(\mathrm{d}x_{1})\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}x_{0}|x_{t}),$
$\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}x_{0}|x_{t})=\frac{\pi_{0}(\mathrm{d}x_{0})\,\mathrm{N}(x_{t};\alpha_{t}x_{0}+\sigma_{t}\mu,\sigma_{t}^{2}\Sigma)}{\pi_{t}(x_{t})}.$ (53)
Indeed, for any bounded measurable function $f$, we get
$\int f(x_{0},x_{t},x_{1})\,\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})\pi_{t}(\mathrm{d}x_{t})$ $=\int f(x_{0},x_{t},x_{1})\,\updelta_{\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}}}(\mathrm{d}x_{1})\pi_{0}(\mathrm{d}x_{0})\mathrm{N}(x_{t};\alpha_{t}x_{0}+\sigma_{t}\mu,\sigma_{t}^{2}\Sigma)\mathrm{d}x_{t}$
$=\int f(x_{0},x_{t},\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}})\,\pi_{0}(\mathrm{d}x_{0})\mathrm{N}(x_{t};\alpha_{t}x_{0}+\sigma_{t}\mu,\sigma_{t}^{2}\Sigma)\mathrm{d}x_{t}$
$=\int f(x_{0},\alpha_{t}x_{0}+\sigma_{t}\mu+\sigma_{t}\Sigma^{1/2}z,\mu+\Sigma^{1/2}z)\,\pi_{0}(\mathrm{d}x_{0})\mathrm{N}(z;0,\mathbf{I}_{d})\mathrm{d}z$
$=\int f(x_{0},\alpha_{t}x_{0}+\sigma_{t}x_{1},x_{1})\,\pi_{0}(\mathrm{d}x_{0})\mathrm{N}(x_{1};\mu,\Sigma)\mathrm{d}x_{1}$
$=\int f(x_{0},x_{t},x_{1})\,\updelta_{\alpha_{t}x_{0}+\sigma_{t}x_{1}}(\mathrm{d}x_{t})\pi_{0}(\mathrm{d}x_{0})\mathrm{N}(x_{1};\mu,\Sigma)\mathrm{d}x_{1}$
@@
$\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})\pi_{t}(\mathrm{d}x_{t})=\updelta_{\alpha_{t}x_{0}+\sigma_{t}x_{1}}(\mathrm{d}x_{t})\pi_{0}(\mathrm{d}x_{0})\pi_{1}(\mathrm{d}x_{1})\;,$ (54)
where the r.h.s. is the joint distribution of the random variables $(X_{0},X_{t},X_{1})$ defined by (48). It then follows that
@@
$=\int q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{s}|x_{0},x_{1})\,\pi_{0}(\mathrm{d}x_{0})\pi_{1}(\mathrm{d}x_{1})$
$=\pi_{s}(\mathrm{d}x_{s})$
@@
$=\int\underbrace{q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{s}|x_{0},\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}})}_{q^{\eta}_{\smash{s|0,t}}(\cdot|x_{0},x_{t})}\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}x_{0}|x_{t})$
where the defined $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(\cdot|x_{0},x_{t})$, up to the notation, is exactly the DDIM bridge transition Song et al. (2021, Equation 7) when $\mu=0_{d}$ and $\Sigma=\mathbf{I}_{d}$. Finally, the Gaussian approximation $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(\cdot|\hat{x}_{0}(x_{t},t),x_{t})$ used at inference, with $\hat{x}_{0}(x_{t},t)\coloneqq\int x_{0}\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\mathrm{d}x_{0}$, is the one solving
$\operatorname*{argmin}_{r_{s|t}(\cdot|x_{t})\in\mathcal{G}_{\eta^{2}_{s}\Sigma}}\mathrm{KL}(\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x_{t})\|r_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x_{t}))\;,$